OpenStudy (anonymous):

Identify which table shows a direct variation. Table 1 x y 2 2 6 12 9 18 Table 2 x y 2 14 6 42 9 63 A. Table 1 B. Table 2 C. Table 1 and Table 2 D. Neither table

OpenStudy (anonymous):

@ganeshie8

OpenStudy (xguardians):

I believe it will be table two. 2/14 = 7 6/42 = 7 9/63 = 7

OpenStudy (xguardians):

Table one doesn't work since 2/2 = 1 6/12 = 2 9/18 = 2

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

@xGuardians

OpenStudy (xguardians):

B.